OpenStudy (anonymous):

Question on statistics: 1. What does variance and standard deviation tell us? 2. The formula for sample variance, why is it as such?? what does Xi mean?

OpenStudy (anonymous):

\[\sum_{}^{?}(X i -Xbar)^2/n\]

OpenStudy (anonymous):

How much time have u got? And how much do u know about probability?

OpenStudy (anonymous):

First, SD^2 = Var(x) = <(x-<x>)^2> where <x> is the mean (by this, I mean the average, x bar = 1/N sum from i=1 to N of x_i, but where N tends to infinity, for small N, average is sufficient.) x is a so-called "random variable" (which can be x_i, a set of numbers for small N). Both SD and Var are measures of how "spread out" a data set is, the spread of a distribution about it's mean or average value.

OpenStudy (anonymous):

what is their difference?

OpenStudy (anonymous):

SD^2 = Var(x)

OpenStudy (anonymous):

both are spread with reference to the sample mean? or population mean?

OpenStudy (anonymous):

It will be better to speak about either a sample (small N) or a population to save specifying everything twice.

OpenStudy (anonymous):

hmmmmm so whats the use of knowing variance and sd

OpenStudy (anonymous):

There are many "statistics", these are just 2 of them, u can think of SD as giving more fine grained information about the sample.

OpenStudy (anonymous):

more detailed you mean?

OpenStudy (anonymous):

Just more....

OpenStudy (anonymous):

hmmm alright.. can you talk to me about estimation of population parameters?

OpenStudy (anonymous):

The simplest way to see it is to take some small samples and then play about calculating both SD and mean (Wolfram does it well, sample and population).

OpenStudy (anonymous):

"estimation of population parameters" U mean from samples?

OpenStudy (anonymous):

i don't know im very confused right now with all the symbols of summation and x here and there

OpenStudy (anonymous):

I agree all the terminology is something of a problem, nothing for it but to get grips with it.... Estimating a parameter just means taking a (hopefully) representative sample to get a statistic.

OpenStudy (anonymous):

find the unbiased estimate of the population mean and population variance from which each of the following samples is drawn:

OpenStudy (anonymous):

unbiased estimate of the population means what

OpenStudy (anonymous):

Well, I would take that to mean that the sample is representative of the population.

OpenStudy (anonymous):

ok..........

OpenStudy (anonymous):

then they give us a lot weird things to calculate

OpenStudy (anonymous):

I don't like calculating things...:-)

OpenStudy (anonymous):

hmmmmm me too

OpenStudy (anonymous):

Like I said, that's a job for software (once u understand the principles)

OpenStudy (anonymous):

ok say there's this summation of (x-100) =52 summation of (x-100)^2 =93 , n=100 how to calculate pop mean and pop variance

OpenStudy (anonymous):

i have to do manually..=(

OpenStudy (anonymous):

I suggest u post that as a new question (it's not that hard).

OpenStudy (anonymous):

new question on the board?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Variance http://en.wikipedia.org/wiki/Standard_deviation There are examples of things here u can look at...

OpenStudy (anonymous):

i need to memorize all the formula to solve this

OpenStudy (anonymous):

sum from i= 1 to n just means let i take the value 1,2,3,.....n

OpenStudy (anonymous):

Any way put up a problem on the board, see if someone will help...

OpenStudy (anonymous):

ok